OpenStudy (anonymous):

If angle E measures (7x -12)º, what is the measure of its complement?

OpenStudy (anonymous):

an angle and it's complement add up to 90 degrees so the compliment of angle E = 90-(7x-12) =90-7x+12 =102-7x